# Working Backwards: Learning to Place by Picking

### 04 Dec 2023 v1

@@ IV-A LPP Data Collection Robustness
Figure 6: Visualization of edges in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
To determine the position of an edge, we fit an ellipse to the detected contact region. If an edge is detected, we calculate the difference between the ideal and detected depth of the edge and use this to calculate the commanded change in EE pose. An example of this is shown in Figure 4. While not all graspable objects contain edges, a large subset of items do as shown in Figure 6.
@@
We conduct an ablation study on CCG and TR where we run self-supervised data collection for a total of 128 episodes and record the number of failures. We plot the results in Figure 5. The naive approach without CCG and without TR fails a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed. The amount of failures drop to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention. Finally, with both CCG and TR, we are able to successfully collect 128 episodes without any human intervention.
@@ IV-B Noise Augmentation Ablation
We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the action allows the policy to better capture this by not having to average over multiple possible expert actions [38]. We also qualitatively observe that policies trained with noise-augmented datasets tend to be better at re-adjusting the object with respect to the goal.
@@ IV-C Comparison to Kinesthetic Teaching
Figure 7: Comparison of success rates of policies trained using a dataset collected by LPP and a traditional kinesthetic teaching approach. We test models trained on an increasing amount of demonstration data. Policies trained on data collected by LPP outperform those trained on data collected from kinesthetic teaching in almost all dataset sizes. The shaded region consists of one standard deviation.
We compare the performance of policies trained using demonstration data from LPP and from kinesthetic teaching. In particular, we investigate whether the quality of the demonstrations differ. To compare, we train two separate policies, both sharing the architecture outlined in Section III-C, on two separate datasets: one collected with LPP and another collected with kinesthetic teaching [39]. We use a Gaussian mixture model as the action representation of both policies. For LPP, we use the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
@@
We hypothesize that the kinesthetic policy performs more poorly than the LPP policy because the overall quality of the human demonstrations is lower than the programmatic or machine-generated demonstrations produced by LPP. Previous work has studied the difficulties and challenges of training policies using human demonstrations when compared to programmatic or machine-generated demonstrations [40, 41, 42]. In particular, human demonstrations can be sub-optimal due to human mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think).
@@ V Limitations and Future Work
The quality and diversity of the collected grasps in LPP are partly determined by the viewpoint of the camera. In this work, we empirically select a viewpoint that is adequate for our use-case. It would be interesting to allow the manipulator to explore and find vantage points for better grasp generation. Furthermore, while tactile sensing is used to improved the robustness of data collection, it could also be used as an extra modality for the policy.
@@ VI Conclusion
We presented LPP, a self-supervised data collection method to collect expert demonstrations for placement tasks. LPP leverages recent advancement in grasp planners and the symmetry between picking and placing to autonomously collect placing demonstrations. LPP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing. We showed that the quality of the demonstrations produced by LPP surpasses that of human demonstrations collected by kinesthetic teaching. LPP provides a promising and simple direction for real world robotic placing demonstration data collection with minimal human labour.

### 20 Mar 2024 v2

@@ III-B Compliant Control for Grasping
Figure 4: Visualization of contact surfaces in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
III-C Tactile Regrasping
Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip significantly once grasped. In the real world, the accumulation of errors from various sources (e.g., camera calibration, controller, learned grasp planner, and noise from the RGB-D sensor and manipulator encoders) results in grasps that are unstable (e.g., a shallow grasp) and not ideal for PvP. To mitigate this issue, we use tactile sensing (specifically, a visuotactile Finger-STS [32] sensor) to preemptively detect if a grasp is stable and perform a regrasp if needed. The Finger-STS is capable of providing multimodal feedback (i.e., both tactile and visual observations). To this end, we design our regrasping module to detect the contact surface area and calculate the relative change in EE pose required to recover a larger contact surface and, thereby, a more stable grasp. We detect regions of contact based on marker displacements and the RGB image from the Finger-STS. If an impartial contact is detected, we calculate the difference between the ideal and detected contact region and use this to calculate the commanded change in EE pose. We note that a simpler and more general random regrasp strategy could also be used. An example of this is shown in Figure 3. We visualize contact surfaces of various household items in Figure 4.
III-C1 Retrieving
We begin the retrieval phase when a valid, stable grasp is detected. An image of the retrieval procedure is shown in Figure 1. We store the poses of the end effector (EE) as well as the respective timestamps throughout the retrieval process. For our work, we measure joint encoder readings at a rate of 120 $\mathrm{H}\mathrm{z}$. The manipulator is set to be compliant along the rotational axes of the EE only, when commanded to move to the pregrasp pose. This allows the manipulator to better comply with any unexpected environment interactions. Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{\text{clearance}}$ centered around a set fixed pose above the scene. That is, given a set fixed pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$ to recover a clearance pose. We use $\sigma_{tr}$ = 2.5 $\mathrm{cm}$ during data collection.
At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose:
where $\mathbf{T}_{0}=\mathbf{T}_{\text{grasp}}$, $\mathbf{T}_{M}=\mathbf{T}_{\text{clearance}}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0. To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a sparser trajectory $\bar{\tau}^{\text{expert}}_{r}$. We extract a total of $\bar{M}+1$ states where $\bar{M}=\lceil\frac{t_{M}-t_{0}}{\Delta t}\rceil$. We do this by finding the nearest pose in terms of time from the dense trajectory $\tau^{\text{expert}}_{r}$ at every $\Delta t$ interval:
$\bar{\tau}^{\text{expert}}_{r}=((\bar{\mathbf{T}}_{0},0),(\bar{\mathbf{T}}_{1},\Delta t),...,(\bar{\mathbf{T}}_{\bar{M}},\bar{M}\Delta t)).$ (2)
In this work, our visual policies operate at 5 $\mathrm{H}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
III-C2 Placing by reversing
We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory
$\bar{\tau}^{\text{expert}}_{p}=(\bar{\mathbf{T}}_{\bar{M}},...,\bar{\mathbf{T}}_{0}).$ (3)
$\Delta\mathbf{T}_{i}=\bar{\mathbf{T}}_{\bar{M}-i}^{-1}\bar{\mathbf{T}}_{\bar{M}-i-1},$ (4)
where $i\in\{0,...,\bar{M}-1\}$ and $\Delta\mathbf{T}_{i}\in SE(3)$ is a relative change in EE pose. We define the expert action as $\mathbf{a}_{i}=(\Delta\mathbf{T}_{i},a_{\text{gripper},i})$, where $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close). To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}-1}$, and store the sequences of observations and actions as an expert trajectory. We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$. We downsample the images to a resolution of $128\times 128$ and also stack the three previous RGB frames, which results in $\mathbf{o}_{i}\in\mathbb{R}^{128\times 128\times 12}$. During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}-1\}$). Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper (i.e. $\Delta\mathbf{T}_{i}=\boldsymbol{\mathbf{I}}$ and $a_{\text{gripper},i}=0$) for N time steps. We use $N=5$, which gives the robot an adequate amount of time to open the gripper and release the object. A single expert demonstration or trajectory is then defined as a set of training tuples
$\tau^{\text{expert}}=\{\mathbf{(o}_{t},\mathbf{a}_{t},\mathbf{o}_{t+1})\}_{t=0}^{T},$ (5)
with a total length of $T=\bar{M}+N$.
III-D Noise-Augmented Data Collection
To increase the coverage of our demonstrations and train more robust policies, we perturb the first $75\%$ of the poses with noise. We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [33], and actions [34]. We also draw inspiration from robust control theory, where injecting isotropic Gaussian noise has been shown to achieve persistent excitation [35], a condition where the training data is informative enough to learn a model that is robust to compounding errors during deployment. We call this variant noise-augmented data collection. In practice, we represent pose commands $\Delta\mathbf{T}$ as a combination of a translation vector $\mathbf{t}\in\mathbb{R}^{3}$ and a rotation vector $\boldsymbol{\theta}\in\mathbb{R}^{3}$. The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis. We perturb the translation of the pose with
$\mathbf{t}_{\text{perturbed}}=\mathbf{t}+\delta\mathbf{t},$ (6)
where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation with
$\boldsymbol{\theta}_{\text{perturbed}}=(\theta+\delta\theta)(\mathbf{e}+\delta\mathbf{e}),$ (7)
where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta\in\mathbb{R}$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$. We perturb both the rotation axis and the rotation angle. We find $\sigma_{t}=0.5~{}$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5~{}$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values. We evaluate the effects of this augmentation on policy performance in Section IV-B
III-E Policy Learning
We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}\,|\,\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by PvP. We use behavioral cloning (BC) [36] to train our policy with a likelihood-based loss or objective function. Given a dataset of $N_{train}$ demonstration trajectories, our loss function is then:
The main backbone of our policy network consists of a convolutional neural network (CNN) based on the ResNet18 architecture [37], followed by a multilayer perceptron (MLP) layer that maps the CNN output into the parameters of the distribution of actions. In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}|\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians. In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [38] and [39], by setting the standard deviation of the Gaussian distributions to be an arbitrarily small value of $61e-$ 6. Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss.
IV Experiments
[…]
Figure 5: Sequence of images from roll outs of place policies trained using data collected with PvP. The policies are able to place objects of varying properties in the scene using images from the wrist camera directly.
We used PvP to collect data and train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 5. In both cases, the place policy had to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle). In this section, we present our quantitative results for the more challenging dishrack loading task as it better represents the type of task, with additional contact constraints from the environment, that PvP was designed for. We study the robustness of PvP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the relative quality of policies trained with data from PvP and from kinesthetic teaching in Section IV-C.
IV-A PvP Data Collection Robustness
Figure 6: Ablation study on compliant control for grasping (CCG) and tactile regrasping (TR). We measured the average number of episodes that could be collected autonomously based on the number of failures. Both CCG and TR play a crucial role in reducing the number of failures to zero, which allowed us to collect the necessary amount of data seamlessly.
We require a robust data collection loop to collect a large number of place demonstrations autonomously with PvP. We studied the effects of compliant control for grasping (CCG) and tactile regrasping (TR), as introduced in Section III, on the robustness of PvP.
We observed two main causes, initially discussed in Section III-A2, that led to misplaced plates during data collection. First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the dishrack. During retrieval, significant object motion occurred at the moment when the object’s contact force with the environment was released. This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the dishrack slot. With PvP, we mitigate this failure mode with CCG. The second failure mode involved unstable grasps that lead to a large relative motion between the object and the gripper. With PvP, we mitigate unstable grasps using TR to preemptively detect unstable grasps and to regrasp if needed.
@@ IV-A PvP Data Collection Robustness
We conducted an ablation study on CCG and TR where we ran self-supervised data collection for a total of 128 episodes and recorded the number of failures. We visualize the results in Figure 6. The naive approach without CCG and without TR failed a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed. The amount of failures dropped to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention. Finally, with both CCG and TR, we were able to successfully collect 128 episodes without any human intervention.
@@ IV-B Noise Augmentation Ablation
We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the action allows the policy to better capture this by not having to average over multiple possible expert actions [40]. We also qualitatively observed that policies trained with noise-augmented datasets were better at re-adjusting the object with respect to the goal.
@@ IV-C Comparison to Kinesthetic Teaching
Figure 7: Comparison of success rates of policies trained using a dataset collected by PvP and a traditional kinesthetic teaching approach. We tested models trained on an increasing amount of demonstration data. Policies trained on data collected by PvP outperform those trained on data collected from kinesthetic teaching for almost all dataset sizes. The shaded region consists of one standard deviation.
We compared the performance of policies trained using demonstration data from PvP and from kinesthetic teaching. In particular, we investigated whether the quality of the demonstrations differ. To compare, we trained two separate policies, both sharing the architecture outlined in Section III-E, on two separate datasets: one collected with PvP and another collected with kinesthetic teaching [41]. We used a Gaussian mixture model as the action representation of both policies. For PvP, we used the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
@@
We hypothesize that the kinesthetic policy performed more poorly than the PvP policy because the overall quality of the human demonstrations was lower than the programmatic or machine-generated demonstrations produced by PvP. Previous work has studied the difficulties and challenges of training policies using human demonstrations when compared to programmatic or machine-generated demonstrations [42, 43, 39]. In particular, human demonstrations can be sub-optimal due to mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think). The authors of [39] note that the relative average trajectory length is a good proxy for the quality of the dataset. The average number of time steps for each demonstration was 29.41 (2.22) for the PvP dataset and 41.66 (5.69) for the kinesthetic dataset. The numbers in the brackets denote variations of one standard deviation. Kinesthetic demonstrations were on average longer and had greater variance in length. In addition, PvP benefits from systematic noise-augmentation, as covered in Section III-D, which a human demonstrator cannot emulate consistently.
@@ V Limitations and Future Work
The quantity, quality and diversity of the collected grasps in PvP are partly determined by the viewpoint of the camera. In this work, we empirically select a viewpoint that is adequate for our use-case. It would be interesting to allow the manipulator to explore and find better grasp poses. Furthermore, while tactile sensing is used to improve the robustness of data collection, it could also be used as an extra modality for the policy. We have introduced a framework for data collection and tested it with two real-world tasks. It would be interesting to test PvP on a larger set of tasks. Lastly, PvP is a language-driven data collection method, and it would be natural to integrate it into a high-level language based planner [9].
@@ VI Conclusion
We presented PvP, a self-supervised data collection method to collect expert demonstrations for placement tasks. PvP leverages recent advancements in grasp planners, tactile sensing, compliant control and the symmetry between picking and placing to autonomously collect placing demonstrations. PvP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing. We showed that the quality of the demonstrations produced by PvP surpasses that of human demonstrations collected by kinesthetic teaching. PvP provides a promising and pragmatic direction for real world robotic placing demonstration data collection with minimal human labour.